OpenStudy (anonymous):

The temperature of the water in a cup falls after the ice has been added. Specific heat capacity of water = 4200 J kg-1 K-1. (i) Calculate the mass of the water in the cup if the lowest temperature reached by the water is 8 degrees C.of the ice is The energy to raise the temperature of the ice to 0 degrees C is 55(4) Joules

OpenStudy (anonymous):

what is the start temperature of water?

OpenStudy (anonymous):

The temperature of the water is 22 degrees C. To change the ice to water without an increase in temperature = 7260 J

OpenStudy (anonymous):

ok just to sum it up: start temperature 22°C end temperature 8 °C cp = 4,2 kJ kg-1K-1 DH of melting = 7,26 kJ

OpenStudy (anonymous):

Yes. Also to to raise the temperature of the ice from -12 Degrees C to 0 Degrees C it takes 55(4) J

OpenStudy (anonymous):

what is 55(4) i mean what is that (4)

OpenStudy (anonymous):

Sorry. It should be 554.4 J

OpenStudy (anonymous):

ok now it makes sense

OpenStudy (anonymous):

just a sec to write it all

OpenStudy (anonymous):

so you have ice at temperature -12 °C which is heated to 0 °C then you have water at 22°C which is cooled to 8°C and you have one phase transition ice to water so lets start by calculating rise of temperature of ice \[Q = m cp \Delta T\] cp(ice) = 2,11 kJ kg-1 K-1 m= Q/cp DT = 554,4 J / 2,11 kJ K-1kg-1 * 12 K m= 0,02189 kg = 21,89 g now lets see what is the situation with liquid water: m= Q/cpDT = 7260 J + 554,4J / 4200 J kg-1K-1 * 14K m= 0,1328 kg = 132,8 g and sumed up 154,69 g of water now that is what i think your answer is

OpenStudy (anonymous):

Convert the mass into Kg.

OpenStudy (anonymous):

i think i made misstake

OpenStudy (anonymous):

could you please write again complete problem in one peace cause i have to catch data all over the page...

OpenStudy (anonymous):

Ok.This question is in two parts. Ice of mass 22g at temperature -12Degrees C is taken from a freezer and placed in a polystyrene cup containing water at a temperature 22 degrees C. (a) Calculate the quantity of heat needed (i) to raise the temperature of the ice from -12 degrees C to 0 degrees C (ii) to change the ice to water without an increase in temperature.

OpenStudy (anonymous):

Then following on:The temperature of the water in a cup falls after the ice has been added. Specific heat capacity of water = 4200 J kg-1 K-1. (i) Calculate the mass of the water in the cup if the lowest temperature reached by the water is 8 degrees C.

OpenStudy (anonymous):

ok and you need to be given heat capacity of ice and energy of phase transition

OpenStudy (anonymous):

are you given heat capacity of ice and energy of phase transition?

OpenStudy (anonymous):

Heat capacity of ice is 2100J kg-1 k-1. Specific latent heat of fusion of ice = 3.3 * 10^5 J kg-1.

OpenStudy (anonymous):

ok now give me a minute to calculate it on paper and then ill write you all here

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

so first lets calculate heat that is needed for ice to be heated to 0°C Q= mcpdT = 22*10-3 kg * 2100 J kg-1 K-1 * 12K = 554,4 J that is answer to (i) then lets calculate heat that is needed to turn ice to water: Q = m * DH = 22*10^-3 kg * 3,3 * 10^5 J kg-1 = 7260 J that is answer to (ii) then lets calculate heat needed for water to be heated from 0 to 8 °C cause that sumed up heat is equal to heat of cooling water from 22 to 8 °C Q= cp m dT = 4200 J kg-1 K-1 * 22*10^-3 kg * 8 K = 739,2 J now lets sum up the heat: Q = 739,2 + 7260 + 554,4 = 8553,6 J and mass of water in cup is: m= Q/ cp dT = 8553,6 J / 4200 J kg-1 K-1 * 14K = 0,1454 kg

OpenStudy (anonymous):

Yes that is the correct answer!

OpenStudy (anonymous):

im a bit rusty but my brain still works :)